OpenStudy (anonymous):

f(x)=x^2-3 why is this an "even" function ?

OpenStudy (raden):

if f(x) = f(-x) then it is an even funtion

OpenStudy (raden):

f(x)=x^2-3 f(-x) = (-x)^2 - 3 = x^2 - 3 = f(x)

OpenStudy (anonymous):

tnks